OpenStudy (anonymous):

find dr/dx for r= tanxsecx What is the rule to solve this problem? How do you solve it?

OpenStudy (anonymous):

I go the answer sec^3xtanx

OpenStudy (anonymous):

Ahh...you would be right...lol

OpenStudy (anonymous):

wanted to know how to change the answer to (sin^2(x)+1) sec^3(x)

OpenStudy (anonymous):

Product rule. Write down tanx, differentiate secx + write down secx, differentiate tanx . Hope this helps!

OpenStudy (loser66):

\[ r'= (tanx)' sec x + (sec x)' tan x\\\huge= sec^3 x + sec^2x tanx = sec^3x ( 1 + \frac{tan x}{secx})\] and \(sec x =\dfrac {1}{cos x}\rightarrow \dfrac{1}{sec x}= cos x\) replace to above \[\huge sec^3 x(1 + tanx cox x)= sec^3 x (1 + sinx)\] cannot = \(sec^3 x ( 1 + sin^2x)\)